Given the demand function P = 20 – 5Q, find the price elasticity of demand when price of the commodity is 5 Birr per unit. Mention if the demand is price elastic or inelastic at this point


1
Expert's answer
2022-02-07T08:42:01-0500

Solution:

Derive inverse demand function:

P = 20 – 5Q

Q = 4 – 0.2P

Q = 4 – 0.2(5) = 4 – 1 = 3

Q = 3

Price elasticity of demand (PED) = "\frac{\triangle Q}{\triangle P} \times \frac{P}{Q}"

"\frac{\triangle Q}{\triangle P}" = -0.2


PED = "-0.2\times \frac{5}{3} = -0.33"

Price elasticity of demand (PED) = 0.33

 

The demand is price inelastic since it less than one. This means that a price change causes a smaller percentage change quantity demanded.